OpenStudy (anonymous):

5x+2y=−5 −6x+15y=4 what do these lines represent ?

OpenStudy (anonymous):

are they parallel or perpendicular or neither?

OpenStudy (abmon98):

change both equation into slop-intercept form 2y=-5x-5 y=-5/2x-5/2 15y=6x+4 y=6/15x+4/15 m=-5/2 m=2/5 they are perpendicular